# One-Shot Affordance Grounding of Deformable Objects in Egocentric Organizing Scenes

### v1

@@ Abstract
Deformable object manipulation in robotics presents significant challenges due to uncertainties in component properties, diverse configurations, visual interference, and ambiguous prompts. These factors complicate both perception and control tasks. To address these challenges, we propose a novel method for One-Shot Affordance Grounding of Deformable Objects (OS-AGDO) in egocentric organizing scenes, enabling robots to recognize previously unseen deformable objects with varying colors and shapes using minimal samples. Specifically, we first introduce the Deformable Object Semantic Enhancement Module (DefoSEM), which enhances hierarchical understanding of the internal structure and improves the ability to accurately identify local features, even under conditions of weak component information. Next, we propose the ORB-Enhanced Keypoint Fusion Module (OEKFM), which optimizes feature extraction of key components by leveraging geometric constraints and improves adaptability to diversity and visual interference. Additionally, we propose an instance-conditional prompt based on image data and task context, effectively mitigates the issue of region ambiguity caused by prompt words. To validate these methods, we construct a diverse real-world dataset, AGDDO15, which includes $15$ common types of deformable objects and their associated organizational actions. Experimental results demonstrate that our approach significantly outperforms state-of-the-art methods, achieving improvements of $6.2\%$, $3.2\%$, and $2.9\%$ in KLD, SIM, and NSS metrics, respectively, while exhibiting high generalization performance. Source code and benchmark dataset will be publicly available at https://github.com/Dikay1/OS-AGDO.
@@ II-A Manipulation and Perception of Deformable Objects
Deformable object manipulation is a critical and complex task in robotics [11, 12]. Compared to rigid or articulated objects, deformable objects pose challenges, including uncertain states and a wide variety of patterns and materials, making their manipulation more difficult. Current research typically employs reinforcement learning [13, 14, 15] or imitation learning approaches [16, 17]. However, precise execution relies on the accurate perception of the object’s state, including deformation and topological structure. Huang et al. [18] propose a self-supervised loss-based method for occlusion reasoning that reconstructs a cloth mesh and uses a dynamic model for planning. UniGarmentManip [7] enables robots to generalize to garment manipulation tasks by learning topological visual correspondences across garments with minimal human input, but it still relies on manual key point identification.
@@ II-B Affordance Grounding of Deformable Objects
Affordance grounding for deformable objects focuses on identifying interaction regions. Existing methods typically rely on operational points, including vertex, skeleton, or edge-based localization using object features [14, 20], motion capture systems combined with human demonstration markers [21], and custom hardware approaches [22, 23]. However, current studies predominantly address single-task scenarios, including folding [8, 24, 25], unfolding [15, 26], grasping [27, 28], and dressing tasks [29, 30]. For instance, Cloth Funnels [31] proposes a heuristic folding operation for garments, setting six key points on a long-sleeve T-shirt. FabricFolding [19] defines four key points for towel manipulation, starting from the top-left corner. Chen et al. [32] compare the frequency of eight grasp points on clothes, finding the back collar is the most frequently used in single-arm fling operations. ReKep [25] introduces a vision-based method for optimizing robotic actions, although manual filtering is still required. Gu et al. [15] apply attention mechanisms to assign higher weights to key points. Sundaresan et al. [33] compare feature-matching methods for correspondence, including SIFT, SURF, and ORB, and explore correspondence using dense neural networks.
@@ III OS-AGDO: Proposed Framework
The proposed framework, shown in Fig. 2, comprises a visual encoder, a text encoder, and a transformer decoder. The input image $I_{\text{ego}}$ is passed through the DINOv2 visual encoder and the ORB-Enhanced Keypoint Fusion Module (OEKFM). DINOv2 extracts dense patch embeddings, producing visual features $F_{v}\in\mathbb{R}^{L\times C}$ and a global [CLS] token. The visual features $F_{v}$ are passed to the Deformable Object Semantic Enhancement Module (DefoSEM), which we propose to enhance functional region features of the target object using dual attention mechanisms , as described in Sec. III-A. The enhanced features are sent to both the transformer decoder and the Meta-Net in the CoCoOp module. The Meta-Net generates dynamic prompts, which, combined with affordance labels, are processed by the CLIP text encoder to produce text embeddings $F_{t}\in\mathbb{R}^{N\times C}$, as explained in Sec. III-C. Simultaneously, we propose OEKFM, which extracts keypoints and generates geometric feature regions using a Gaussian kernel, as detailed in Sec. III-B. The transformer decoder receives the visual features $F_{v}$, text embeddings $F_{t}$, and the [CLS] token, producing an initial affordance prediction. This prediction $F_{\text{pred}}$ is subsequently optimized using the geometric features from OEKFM.
@@ III-A DefoSEM
As shown in the orange box in Fig. 2, we enhance the visual features extracted by DINOv2. The DefoSEM combines image label $L$ with visual features $F_{v}$, improving feature learning through a multi-scale attention mechanism.
First, adaptive average pooling is applied to downsample the image labels, reducing the spatial dimensions while preserving global information. Next, a convolutional operation is used to align the channel dimensions of the label features with the visual features. The visual features are then adjusted to generate $\tilde{F_{v}}$, and spatial attention weights $W_{s}$ are computed. These weights modulate the visual features along the spatial dimension, emphasizing important spatial regions. Simultaneously, the adjusted visual features $\tilde{F_{v}}$ and the dimension-processed label features $\tilde{L}$ are concatenated along the channel dimension. A convolution operation then generates channel attention weights $W_{c}$, which are used to weigh the visual features and highlight important channel-specific information. Finally, the adjusted visual features $\tilde{F_{v}}$, channel attention weights $W_{c}$, and spatial attention weights $W_{s}$ are fused to generate the enhanced features $F_{v}^{{}^{\prime}}$. The fusion formula is as follows:
$F_{v}^{{}^{\prime}}=\tilde{F_{v}}~{}\times(0.5+W_{c})\times(0.5+W_{s}).$
@@ III-B ORB-Enhanced Keypoint Fusion Module
As shown in the yellow box in Fig. 2, we introduce a new branch after inputting the image $I_{\text{ego}}$, referred to as OEKFM. First, we apply feature matching using the ORB algorithm to extract keypoints of the deformable object. These keypoints are then processed using a Gaussian kernel to generate a feature region $M$.
@@
$P=F_{\text{pred}}\odot M^{\prime}.$
@@
$M^{\prime}=1+\frac{M-M_{\text{min}}}{M_{\text{max}}-M_{\text{min}}+\epsilon}.$
@@
$P_{\text{final}}=\text{clamp}(P,0,1).$
@@
$\text{clamp}(x,0,1)=\begin{cases}0,&\text{if }x<0,\\ x,&\text{if }0\leq x\leq 1,\\ 1,&\text{if }x>1.\end{cases}$
@@ III-C Instance-conditional Prompt
As illustrated in the purple module of Fig. 2, $v_{m}$ represents the learnable context vector for the $m$-th input (where $m{\in}\{1,2,\dots,M\}$), $x$ denotes the feature vector of the image, and $\pi$ represents the instance-conditional generated by Meta-Net. By incorporating the CoCoOp method, $N$ neural networks are constructed to generate $N$ context tokens. Specifically, a lightweight neural network, called Meta-Net, is learned on top of the context vector. This network generates a conditional token for each input, which is combined with the task-specific affordance labels. The current context prompt is given by: $v_{m}(x){=}v_{m}{+}\pi$. The initialized learnable context vectors $\{v_{1}(x),v_{2}(x),\dots,v_{M}(x)\}$ are inserted before the text CLS token, allowing for shared functionality across all. Thus, the text prompt is: $t_{i}(x){=}\{v_{1}(x),v_{2}(x),\dots,v_{M}(x),c_{i}\}.$ The final prediction is calculated as:
$p(y|x)=\frac{exp(sim(x,g(t_{y}(x)))/\tau)}{\sum_{i=1}^{K}exp(sim(x,g(t_{y}(x)))/\tau)}.$
@@ IV-A Dataset Collection
Currently, datasets for one-shot learning in deformable object affordance grounding in organizing scenes are scarce. To address this, we have constructed a dataset using egocentric images from diverse, high-quality sources. We selected egocentric images from diverse, high-quality sources to ensure the dataset’s completeness. Portions of the clothes and towel data come from DeepFashion  [37] and the FabricFolding Dataset [19]. To prevent inadvertently recognizing deformable objects, such as operators’ clothes, during detection in exocentric views, only egocentric images were included. To align with manipulation tasks. Consequently, images from DeepFashion containing human figures or body parts were excluded. A random subset of images from the FabricFolding Dataset was chosen for diversity. Additional data for other deformable objects, along with extra clothes and towel categories, were sourced from freely licensed websites. All deformable objects in the dataset have a coverage rate of over $65\%$. The final dataset contains $451$ images, with $122$ from existing datasets and $329$ from public sources.
@@ V-B Results of Functional Affordance Grounding
TABLE I: Comparison to state-of-the-art methods on the AGDDO15 dataset for relevant tasks (↑/↓ indicates higher/lower is better)
@@ V-C Ablation Study
Number of keypoint features. As shown in Table III, we evaluate the effect of different feature quantities, $n{=}\{300,400,500,1000\}$, on the KLD, SIM, and NSS metrics. The results indicate that the optimal performance across all metrics is achieved when $n{=}400$. This finding aligns with our algorithm design principle: an excessively high number of features may cause the model to focus on background regions or other irrelevant areas, while an insufficient number of features can result in inadequate extraction of critical information, such as edge details.
@@ V-D Performance on Real Deformable Objects
As shown in Fig. 7, we present the results of our method on deformable objects. The left panel illustrates the affordance-based organizing strategy. where the model sequentially identifies predefined functional regions based on prompts. For instance, when organizing a short-sleeve T-shirt, the model first identifies the “grasp_sleeve” region, followed by the “put_center” region, then “grasp_shoulder,” and finally “put_hem.” The first row displays a visualization of our method’s affordance grounding. After identifying all affordance regions, we proceed with the robotic arm operation experiments. In the experiment involving shorts, our annotation design specified “pick” actions in the two right-side regions and “place” actions in the two left-side regions (see “Plan A”). However, after learning from the top-to-bottom annotations in the towel and pants categories, a new strategy emerged for shorts, following a top-to-bottom pattern in which “pick” actions occurred at the two waist regions, and “place” actions were performed at the two pant leg regions. Notably, in the experiment, this strategy successfully completed the organization task (see “Plan B”). These results demonstrate that our method effectively facilitates affordance grounding for deformable objects and validates the practical utility of the model’s learned strategy.
Figure 6: The real-world experimental platform for deformable objects
Figure 7: Examples of deformable objects in organizing scenes based on our OS-AGDO.

### v2

@@ Abstract
Deformable object manipulation in robotics presents significant challenges due to uncertainties in component properties, diverse configurations, visual interference, and ambiguous prompts. These factors complicate both perception and control tasks. To address these challenges, we propose a novel method for One-Shot Affordance Grounding of Deformable Objects (OS-AGDO) in egocentric organizing scenes, enabling robots to recognize previously unseen deformable objects with varying colors and shapes using minimal samples. Specifically, we first introduce the Deformable Object Semantic Enhancement Module (DefoSEM), which enhances hierarchical understanding of the internal structure and improves the ability to accurately identify local features, even under conditions of weak component information. Next, we propose the ORB-Enhanced Keypoint Fusion Module (OEKFM), which optimizes feature extraction of key components by leveraging geometric constraints and improves adaptability to diversity and visual interference. Additionally, we propose an instance-conditional prompt based on image data and task context, which effectively mitigates the issue of region ambiguity caused by prompt words. To validate these methods, we construct a diverse real-world dataset, AGDDO15, which includes $15$ common types of deformable objects and their associated organizational actions. Experimental results demonstrate that our approach significantly outperforms state-of-the-art methods, achieving improvements of $6.2\%$, $3.2\%$, and $2.9\%$ in KLD, SIM, and NSS metrics, respectively, while exhibiting high generalization performance. Source code and benchmark dataset are made publicly available at https://github.com/Dikay1/OS-AGDO.
@@ II-A Manipulation and Perception of Deformable Objects
Deformable object manipulation is a critical and complex task in robotics [11, 12]. Compared to rigid or articulated objects, deformable objects pose challenges, including uncertain states and diverse patterns and materials, making their manipulation more difficult. Current research typically employs reinforcement learning [13, 14, 15] or imitation learning approaches [16, 17]. However, precise execution demands accurate perception of object state (e.g., deformation and topology). Huang et al. [18] propose a self-supervised loss-based method for occlusion reasoning that reconstructs a cloth mesh and uses a dynamic model for planning. UniGarmentManip [7] enables robots to generalize to garment manipulation tasks by learning topological visual correspondences across garments with minimal human input, but it still relies on manual key point identification.
@@ II-B Affordance Grounding of Deformable Objects
Affordance grounding for deformable objects focuses on identifying interaction regions. Existing methods typically rely on operational points, including vertex, skeleton, or edge-based localization using object features [14, 20], motion capture systems combined with human demonstration markers, and custom hardware approaches [21, 22]. However, current studies predominantly address single-task scenarios, including folding [8, 23, 24], unfolding [15, 25], grasping [26, 27], and dressing tasks [28, 29]. For instance, Cloth Funnels [30] proposes a heuristic folding operation for garments, setting six key points on a long-sleeve T-shirt. FabricFolding [19] defines four key points for towel manipulation, starting from the top-left corner. Chen et al. [31] compare the frequency of eight grasp points on clothes, finding the back collar is the most frequently used in single-arm fling operations. ReKep [24] introduces a vision-based method for optimizing robotic actions, although manual filtering is still required. Gu et al. [15] apply attention mechanisms to assign higher weights to key points. Sundaresan et al. [32] compare feature-matching methods for correspondence, including SIFT, SURF, and ORB, and explore correspondence using dense neural networks.
@@ III OS-AGDO: Proposed Framework
The proposed framework, shown in Fig. 2, comprises a visual encoder, a text encoder, and a transformer decoder. The input image $I_{\text{ego}}$ is passed through the DINOv2 [36] visual encoder and the ORB-Enhanced Keypoint Fusion Module (OEKFM). DINOv2 extracts dense patch embeddings, producing visual features $F_{v}\in\mathbb{R}^{L\times C}$ and a global [CLS] token. The visual features $F_{v}$ are passed to the Deformable Object Semantic Enhancement Module (DefoSEM), which we propose to enhance functional region features of the target object using dual attention mechanisms, as described in Sec. III-A. The enhanced features are sent to both the transformer decoder and the Meta-Net in the CoCoOp module. The Meta-Net generates dynamic prompts, which, combined with affordance labels, are processed by the CLIP text encoder to produce text embeddings $F_{t}\in\mathbb{R}^{N\times C}$, as explained in Sec. III-C. Simultaneously, we propose OEKFM, which extracts keypoints and generates geometric feature regions using a Gaussian kernel, as detailed in Sec. III-B. The transformer decoder receives the visual features $F_{v}^{{}^{\prime}}$, text embeddings $F_{t}$, and the [CLS] token, producing an initial affordance prediction. This prediction $F_{\text{pred}}$ is subsequently optimized using the geometric features from OEKFM.
@@ III-A DefoSEM
As shown in the orange box in Fig. 2, we enhance the visual features extracted by DINOv2. The DefoSEM combines image label $L$ with visual features $F_{v}$, improving feature learning through a multi-scale gating mechanism.
First, adaptive average pooling is applied to downsample the image labels, reducing the spatial dimensions while preserving global information. Next, a convolutional operation is used to align the channel dimensions of the label features with the visual features. The visual features are then adjusted to generate $\tilde{F_{v}}$, and spatial gating weights $W_{s}$ are computed. These weights modulate the visual features along the spatial dimension, emphasizing important spatial regions. Simultaneously, the adjusted visual features $\tilde{F_{v}}$ and the dimension-processed label features $\tilde{L}$ are concatenated along the channel dimension. A convolution operation then generates channel gating weights $W_{c}$, which are used to weigh the visual features and highlight important channel-specific information. Finally, the adjusted visual features $\tilde{F_{v}}$, channel gating weights $W_{c}$, and spatial gating weights $W_{s}$ are fused to generate the enhanced features $F_{v}^{{}^{\prime}}$. The fusion formula is as follows:
$F_{v}^{{}^{\prime}}=\tilde{F_{v}}~{}\times(0.5+W_{c})\times(0.5+W_{s}).$ (1)
@@ III-B ORB-Enhanced Keypoint Fusion Module
As shown in the yellow box in Fig. 2, we introduce a new branch after inputting the image $I_{\text{ego}}$, referred to as OEKFM. First, we apply feature matching using the ORB algorithm [37] to extract keypoints of the deformable object. These keypoints are then processed using a Gaussian kernel to generate a feature region $M$.
@@
$P=F_{\text{pred}}\odot M^{\prime}.$ (2)
@@
$M^{\prime}=1+\frac{M-M_{\text{min}}}{M_{\text{max}}-M_{\text{min}}+\epsilon}.$ (3)
@@
$P_{\text{final}}=\text{clamp}(P,0,1).$ (4)
@@
$\text{clamp}(x,0,1)=\begin{cases}0,&\text{if }x<0,\\ x,&\text{if }0\leq x\leq 1,\\ 1,&\text{if }x>1.\end{cases}$ (5)
@@ III-C Instance-conditional Prompt
As illustrated in the purple module of Fig. 2, $v_{m}$ represents the learnable context vector for the $m$-th input (where $m{\in}\{1,2,\dots,M\}$), $x$ denotes the feature vector of the image, and $\pi$ represents the instance-conditional generated by Meta-Net. By incorporating the CoCoOp method [38], $N$ neural networks are constructed to generate $N$ context tokens. Specifically, a lightweight neural network, called Meta-Net, is learned on top of the context vector. This network generates a conditional token for each input, which is combined with the task-specific affordance labels. The vector $v_{m}$ is a static learnable parameter, and $v_{m}(x)$ denotes its instance-conditioned version: $v_{m}(x){=}v_{m}{+}\pi$. The initialized learnable context vectors $\{v_{1}(x),v_{2}(x),\dots,v_{M}(x)\}$ are inserted before the text CLS token, allowing for shared functionality across all. Thus, the text prompt is: $t_{i}(x){=}\{v_{1}(x),v_{2}(x),\dots,v_{M}(x),c_{i}\}$. Here $g(\cdot)$ is the text encoder. The final prediction is calculated as:
$p(y|x)=\frac{exp(sim(x,g(t_{y}(x)))/\tau)}{\sum_{i=1}^{K}exp(sim(x,g(t_{y}(x)))/\tau)}.$ (6)
@@ IV-A Dataset Collection
Existing datasets for one-shot affordance learning of deformable objects in organizing scenes are scarce. To address this, we created a comprehensive dataset of egocentric images sourced from diverse, high-quality origins to ensure its robustness and completeness. Portions of the clothes and towel data come from DeepFashion  [39] and the FabricFolding Dataset [19]. To prevent inadvertently recognizing deformable objects, such as operators’ clothes, during detection in exocentric views, only egocentric images were included to align with manipulation tasks. Consequently, images from DeepFashion containing human figures or body parts were excluded. A random subset of images from the FabricFolding Dataset was chosen for diversity. Additional data for other deformable objects, along with extra clothes and towel categories, were sourced from freely licensed websites. All deformable objects in the dataset have a coverage rate of over $65\%$. The final dataset contains $451$ images, with $122$ from existing datasets and $329$ from public sources.
@@ V-B Results of Functional Affordance Grounding
TABLE I: Comparison to SOTA methods on the AGDDO15 dataset for relevant tasks (↑/↓ indicates higher/lower is better)
@@ V-C Ablation Study
Number of keypoint features. As shown in Table III, we evaluate the effect of different feature quantities, $n{=}\{200,400,600,800,1000\}$, on the KLD, SIM, and NSS metrics. The results indicate that the optimal performance across all metrics is achieved when $n{=}400$. This finding aligns with our algorithm design principle: an excessively high number of features may cause the model to focus on background regions or other irrelevant areas, while an insufficient number of features can result in inadequate extraction of critical information, such as edge details.
@@ V-D Performance on Real Deformable Objects
As shown in Fig. 7, we present the results of our method on deformable objects. The left panel illustrates the affordance-based organizing strategy where the model sequentially identifies predefined functional regions based on prompts. For instance, when organizing a short-sleeve T-shirt, the model first identifies the “grasp_sleeve” region, followed by the “put_center” region, then “grasp_shoulder,” and finally “put_hem.” The first row displays a visualization of our method’s affordance grounding. After identifying all affordance regions, we proceed with the robotic arm operation experiments. In the experiment involving shorts, our annotation design specified “pick” actions in the two right-side regions and “place” actions in the two left-side regions (see “Plan A”). However, after learning from the top-to-bottom annotations in the towel and pants categories, a new strategy emerged for shorts, following a top-to-bottom pattern in which “pick” actions occurred at the two waist regions, and “place” actions were performed at the two pant leg regions. Notably, in the experiment, this strategy successfully completed the organization task (see “Plan B”). These results demonstrate that our method effectively facilitates affordance grounding for deformable objects and validates the practical utility of the model’s learned strategy.
Figure 6: Real-world experimental platform for deformable objects.
Figure 7: Examples of deformable object affordance grounding in organizing scenes based on our OS-AGDO.